# MCDFN: Supply Chain Demand Forecasting via an Explainable Multi-Channel Data Fusion Network Model Integrating CNN, LSTM, and GRU

### v3

@@ 3.2.6 MLP
These equations describe the forward propagation process of an MLP, capturing the transformations from the input through the hidden layers to the output, with each layer applying a linear transformation followed by a non-linear activation function to produce the final prediction.
@@ 4 Results
In the results section, we present the evaluation metrics used in our research, the comparative analysis of the implemented DL models, the generation of final outputs, statistical tests comparing predicted and actual values, statistical significance testing of model performance, and XAI interpretation of MCDFN.
@@ 4.2 Comparative Analysis of Results
Table 2 presents the comparative evaluation of various models, including BiLSTM, CNN, RNN, Stacked LSTM, Vanilla LSTM, MLP, GRU, and Proposed MCDFN. A comparative analysis of these results provides valuable insights into the effectiveness of each model for the given task.
Table 2: Performance evaluation of the benchmarked models on validation and test datasets
@@
Among the models evaluated, the CNN-LSTM-GRU architecture, proposed as MCDFN, demonstrated promising performance (see Figure 10). It achieved the lowest test loss and competitive scores across all evaluation metrics compared to other models. The model exhibited a balanced trade-off between accuracy and complexity, leveraging the strengths of CNNs, BiLSTM, stacked LSTMs, and GRUs. On the other hand, the Vanilla LSTM model, despite being a fundamental architecture, struggled with the test data, yielding the highest test loss and relatively inferior performance in terms of MSE, RMSE, MAE, and MAPE. This may suggest limitations in capturing the temporal dependencies and patterns in the data, leading to suboptimal forecasting accuracy. The BiLSTM, RNN, and GRU models performed comparably, showcasing moderate performance in test evaluation metrics. While these models demonstrated the ability to capture temporal dependencies to some extent, they fell short of achieving the same level of accuracy as the proposed MCDFN architecture. The CNN model exhibited inferior performance compared to other DL architectures, particularly regarding RMSE and MAPE, indicating challenges in effectively leveraging spatial features for demand forecasting tasks. The MLP model, while offering simplicity in architecture, lagged behind in forecasting accuracy compared to more complex DL models, highlighting the importance of leveraging sequential and temporal information for accurate demand predictions.
@@ 4.5 Statistical Significance Testing of Model Performance
The p-value is calculated using the t-distribution to determine the probability of observing the t-statistic under the null hypothesis. The significance level ($\alpha$) is set to 0.05. Table 4 shows the t-values and p-values for the MSE, MAE, and MAPE comparisons between the MCDFN model and various baseline models. The p-values in Table 4 indicate that all comparisons between the MCDFN model and other baseline models are statistically significant at the $\alpha=0.05$ level for all three metrics (MSE, MAE, and MAPE). This suggests that the performance of the MCDFN model is significantly different from the baseline models across all considered metrics. The MCDFN model consistently shows significant improvements over the other models regarding MSE, MAE, and MAPE. Vanilla LSTM and CNN models, in particular, show the highest t-values, indicating large performance differences compared to the MCDFN model.
@@ 4.6 XAI Interpretation
We implemented ShapTime and PFI, both model-agnostic XAI techniques, to elucidate the black-box nature of the MCDFN model.
@@ 4.6.1 ShapTime
Theoretically, if we were to perturb the training data based on the explanation outcomes—for instance, substituting critical super-times with less essential ones—the model’s predictive performance would be markedly deteriorating. In our example, this point is illustrated by replacing the crucial $t_{1}$ with the least influential $t_{6}$ and $t_{8}$ with $t_{3}$. This evaluation approach aligns with the conventional practice in the eXplainable Artificial Intelligence (XAI) domain, commonly called sensitivity analysis.
@@ 4.6.2 Permutation Feature importance
The permutation feature importance (PFI) scores indicate the degree to which each feature influences the predictive performance of our proposed model (see Figure 13). Features with higher importance scores contribute more significantly to the accuracy of the demand forecasts generated by the model. Specifically, the “Week_sin” feature emerges as the most influential predictor, followed by “Month_cos” and “Week_cos,” indicating the pronounced impact of weekly and monthly cyclical patterns on demand fluctuations. Moreover, the temporal features related to year and month exhibit substantial importance, underscoring the seasonal trends inherent in demand dynamics.
@@ 5 Discussion
Demand forecasting in this study was performed at an aggregate product family level. Our results confirm the importance of macro-economic factors in aggregate demand prediction, with the MCDFN model effectively capturing and incorporating complex, non-linear relationships among various predictive variables. Accurate aggregate forecasts can be translated into precise demand predictions at the SKU level, potentially reducing inefficiency and the bullwhip effect prevalent in non-seasonal demand industries (Cachon et al.,, 2007). Effective demand forecasting can significantly enhance SC efficiency, yet eliminating the bullwhip effect requires a system-wide perspective and improved collaboration among SC participants (Gaur et al.,, 2005). While SC integration remains challenging, leveraging BDA capabilities can yield favorable outcomes, especially for firms further removed from the end consumer. This research provides valuable insights into this aspect. Despite the sophisticated analytical methods, many firms still rely on judgmental adjustments in demand forecasting (Boulaksil and Franses,, 2009; Moritz et al.,, 2014; Sanders and Manrodt,, 2003). Our findings support the continued importance of expert judgment in refining forecasts, as highlighted by Sanders and Manrodt, who found a high prevalence of judgmental adjustments among practitioners using model-based methods (Sanders and Manrodt,, 2003).
The power of the MCDFN lies in its ability to integrate different data sources or features as inputs for each channel of the model. Although our current study uses a single data source as input for multiple channels (CNN, LSTM, GRU), this approach can be extended to leverage multiple data modalities. By processing diverse types of data through different channels, the MCDFN can capture a broader range of patterns and dependencies, enhancing the robustness and accuracy of the forecasts. This versatility makes MCDFN a powerful tool for handling complex and multifaceted datasets, providing a significant edge over traditional and single-channel models. To address the “black box” nature of MCDFN, we employ XAI techniques such as ShapTime and PFI. These methods provide insights into the MCDFN’s decision-making processes, helping stakeholders understand and trust the model’s predictions. By elucidating which features contribute most to the forecast, XAI enhances MCDFN’s transparency and interpretability, making it easier for decision-makers to adopt and integrate advanced AI solutions into their operations.
@@ 6 Conclusions
Despite the promising results, our study has several limitations. First, the computational complexity of MCDFN is higher than traditional models, which may limit its applicability in real-time forecasting scenarios or environments with limited computational resources. Second, the performance of MCDFN heavily relies on the quality and quantity of the input data, making it sensitive to noisy or incomplete datasets. Third, while XAI techniques provide some level of interpretability, the model’s complexity can still challenge stakeholders to understand the underlying decision-making processes fully. Fourth, our study considered a single product and a retail shop, potentially introducing selective biases due to geographical positions and other demographic variables.
Future research could address these limitations by exploring several avenues. Conducting a sensitivity analysis of window size to assess its impact on model performance. Implementing advanced data augmentation techniques and robust preprocessing pipelines can enhance the model’s performance on noisy or incomplete datasets. Combining MCDFN with other state-of-the-art techniques, such as attention mechanisms and transformers, could improve its predictive capabilities and interoperability. Comparing performance with generalized DL forecasting methods developed by industry experts, such as N-BEATS (ElementAI) and DeepAR (Amazon). Investigating the scalability of MCDFN in distributed computing environments and deploying it in real-time systems can demonstrate its practical applicability. Extending MCDFN to other domains, such as healthcare or finance, can validate its versatility and generalizability across different types of time series data. Identifying auxiliary variables that affect product demand and studying their influence on forecasting performance.
By addressing these limitations and exploring these future directions, we can further enhance the effectiveness and applicability of MCDFN in various demand forecasting and time series analysis tasks.
@@
Conflicts of Interest
The authors whose names are listed immediately below certify that they have no affiliations with or involvement in any organization or entity with any financial or non-financial interest in the subject matter or materials discussed in this manuscript.

### v5

@@ 3.2.6 FCN
These equations describe the forward propagation process of an FCN, capturing the transformations from the input through the hidden layers to the output, with each layer applying a linear transformation followed by a non-linear activation function to produce the final prediction.
@@ 4 Results
For performance evaluation, our architecture utilized Mean Absolute Error (MAE), Mean Squared Error (MSE), Root Mean Squared Error (RMSE), and Mean Absolute Percentage Error (MAPE). These metrics were chosen due to their widespread use in demand forecasting and their ability to provide complementary insights into model performance. Below, we justify the choice of these metrics and discuss their relevance to SC forecasting.
@@ 4.1.1 Mean Absolute Error (MAE)
MAE is particularly useful for SC forecasting because it provides an intuitive measure of average error magnitude without penalizing larger deviations disproportionately. This makes it suitable for evaluating models where consistent accuracy across all predictions is critical, such as inventory management and order fulfillment.
@@ 4.1.2 Mean Squared Error (MSE)
MSE is sensitive to outliers due to the squaring of errors, making it particularly useful for identifying models that perform poorly on extreme demand spikes or drops. In SC contexts, minimizing MSE ensures that the model avoids large prediction errors, which can lead to stockouts or overstocking.
@@ 4.1.3 Root Mean Squared Error (RMSE)
RMSE is widely used in forecasting because it emphasizes larger errors more than MAE, providing a balanced view of overall model performance. In SC applications, RMSE helps prioritize models that minimize significant deviations, which are often costly in terms of lost sales or excess inventory.
@@ 4.1.4 Mean Absolute Percentage Error (MAPE)
MAPE is particularly valuable in SC forecasting because it provides a relative error measure, making it easier to interpret and compare across different datasets or products. However, it is sensitive to zero or near-zero actual values, which can occur in intermittent demand scenarios.
4.1.5 Theil’s U Statistic
Theil’s U statistic is a relative accuracy measure used to evaluate the performance of a predictive model against a naive benchmark. It is particularly useful in regression and time-series forecasting, where assessing the model’s effectiveness relative to simple heuristics is essential. Theil’s U statistic is computed as follows:
$U=\frac{\sqrt{\frac{1}{n}\sum_{t=1}^{n}(y_{t}-\hat{y}_{t})^{2}}}{\sqrt{\frac{1}{n}\sum_{t=1}^{n}(y_{t}-y_{t-1})^{2}}}$ (25)
where $y_{t}$ represents the actual observed value at time $t$, $\hat{y}_{t}$ is the predicted value at time $t$, $y_{t-1}$ is the previous observed value (naive forecast), $n$ is the total number of observations. The numerator represents the RMSE of the benchmarked model, while the denominator represents the RMSE of a naive forecast, typically assuming that the best prediction for the next step is the previously observed value. $U=1$ indicates that the predictive model performs equivalently to the naive benchmark. $U>1$ suggests that the model performs worse than the naive forecast. $U<1$ signifies that the predictive model outperforms the naive approach, demonstrating better predictive accuracy. $U=0$ represents a perfectly accurate model with no prediction error.
@@ 4.2 Comparative Analysis of Results
Table 2 presents the comparative evaluation of various models, including BiLSTM, CNN, RNN, Stacked LSTM, Vanilla LSTM, FCN, GRU, and Proposed MCDFN. A comparative analysis of these results provides valuable insights into the effectiveness of each model for the given task.
Table 2: Performance evaluation of the benchmarked models on validation and test datasets. Bold indicates the best test results.
@@
Table 3: Ablation study results for one or more modules of MCDFN. Bold indicates the best test results.
Model Loss MSE RMSE MAE MAPE (%)
w/o BiLSTM branch 24.5677 23.7164 5.1299 4.2332 21.0827
w/o CNN branch 24.6199 23.9214 5.5051 4.1633 21.6743
w/o BiGRU branch 24.4913 23.8372 5.0736 4.4519 21.6472
w/o Stacked LSTM branch 25.0792 24.8041 5.1571 4.1067 21.996
w/o BiLSTM and CNN branch 25.1304 24.0481 5.1470 4.148 21.4381
w/o CNN and BiGRU branch 24.5663 24.4179 5.3016 4.0947 21.6999
w/o BiGRU and Stacked LSTM branch 25.1313 24.3707 5.1925 4.0529 21.1524
w/o Stacked LSTM and BiLSTM branch 25.2962 24.5997 5.1558 4.3552 21.2836
w/o CNN and Stacked LSTM branch 24.5836 24.3847 5.3117 4.1446 21.2164
w/o BiLSTM and BIGRU branch 25.4159 24.4595 5.1338 4.0763 21.1825
Proposed MCDFN 23.7287 23.5738 4.8553 3.9991 20.1575
@@
Table 4: Comparative analysis of model performance using Theil’s U statistic, number of parameters, and computational efficiency
Model Theil’s U statistic # Params Training time (ms) Inference time (ms)
BiLSTM 0.1097 657,438 $\approx$7488 $\approx$122
CNN 0.1158 191,006 $\approx$2085 $\approx$940
RNN 0.1051 133,022 $\approx$5148 $\approx$364
Stacked LSTM 0.1331 3,631,134 $\approx$15210 $\approx$124
Vanilla LSTM 0.1528 9,57,150 $\approx$1911 $\approx$74
FCN 0.1184 6145 $\approx$7800 $\approx$386
GRU 0.1082 290,334 $\approx$5440 $\approx$474
Proposed MCDFN 0.1181 1,123,358 $\approx$35100 $\approx$32
Among the models evaluated, the CNN-LSTM-GRU architecture, proposed as MCDFN, demonstrated promising performance (see Figure 10). It achieved the lowest test loss and competitive scores across all evaluation metrics compared to other models. The model exhibited a balanced trade-off between accuracy and complexity, leveraging the strengths of CNNs, BiLSTM, stacked LSTMs, and GRUs. On the other hand, the Vanilla LSTM model, despite being a fundamental architecture, struggled with the test data, yielding the highest test loss and relatively inferior performance in terms of MSE, RMSE, MAE, and MAPE. This may suggest limitations in capturing the temporal dependencies and patterns in the data, leading to suboptimal forecasting accuracy. The BiLSTM, RNN, and GRU models performed comparably, showcasing moderate performance in test evaluation metrics. While these models demonstrated the ability to capture temporal dependencies to some extent, they fell short of achieving the same level of accuracy as the proposed MCDFN architecture. The CNN model exhibited inferior performance compared to other DL architectures, particularly regarding RMSE and MAPE, indicating challenges in effectively leveraging spatial features for demand forecasting tasks. The FCN model, while offering simplicity in architecture, lagged behind in forecasting accuracy compared to more complex DL models, highlighting the importance of leveraging sequential and temporal information for accurate demand predictions.
MCDFN achieves competitive performance with Theil’s U statistic of 0.1181, demonstrating its robustness in predictive accuracy, as shown in Table 4. While it has a relatively higher parameter count (1,123,358) and longer training time ($\approx$35,100 ms), it significantly outperforms all baseline models in inference efficiency with an inference time of just $\approx$32 ms. In contrast, BiLSTM and GRU exhibit slightly lower Theil’s U values (0.1097 and 0.1082, respectively) but require higher inference times ($\approx$122 ms and $\approx$474 ms). Stacked LSTM, despite its deeper architecture, shows the highest Theil’s U statistic (0.1331) along with the longest training time ($\approx$15,210 ms), indicating potential overfitting.
@@
4.3 Ablation Studies
The ablation study presented in Table 3 compares the performance of various model configurations, each missing one or more components (such as BiLSTM, CNN, BiGRU, and Stacked LSTM branches), against the fully configured proposed model, MCDFN. The proposed MCDFN model achieves the lowest values across all metrics, indicating superior performance.
By removing individual or paired branches, performance generally deteriorates across all metrics. For instance, removing the BiLSTM and CNN branches results in a notable increase in RMSE (5.15%) and MAPE (21.44%). Similarly, removing the Stacked LSTM and BiLSTM branches results in one of the highest loss values (25.30%), indicating that both these branches contribute significantly to the model’s accuracy. This analysis highlights the importance of each component within the proposed MCDFN structure, with the complete architecture yielding the most accurate predictions.
@@ 4.6 Statistical Significance Testing of Model Performance
The p-value is calculated using the t-distribution to determine the probability of observing the t-statistic under the null hypothesis. The significance level ($\alpha$) is set to 0.05. Table 6 shows the t-values and p-values for the MSE, MAE, and MAPE comparisons between the MCDFN model and various baseline models. The p-values in Table 6 indicate that all comparisons between the MCDFN model and other baseline models are statistically significant at the $\alpha=0.05$ level for all three metrics (MSE, MAE, and MAPE). This suggests that the performance of the MCDFN model is significantly different from the baseline models across all considered metrics. The MCDFN model consistently shows significant improvements over the other models regarding MSE, MAE, and MAPE. In particular, Vanilla LSTM and CNN models show the highest t-values, indicating large performance differences compared to the MCDFN model.
@@ 4.7 XAI Interpretation
To address the “black box” nature of the MCDFN model, we employed two widely recognized XAI techniques: ShapTime and PFI. These methods provide insights into the model’s decision-making process, enhancing transparency and trust in its predictions.
@@ 4.7.1 ShapTime
Where: - $y_{t}$: Predicted value at time $t$, - $\phi_{t_{i}}$: SHAP value for super-time $t_{i}$, - $\omega_{t}$: Residual error term.
@@
Theoretically, if we were to perturb the training data based on the explanation outcomes—for instance, substituting critical super-times with less essential ones—the model’s predictive performance would be markedly deteriorating. In our example, this point is illustrated by replacing the crucial $t_{1}$ with the least influential $t_{6}$ and $t_{8}$ with $t_{3}$. This evaluation approach aligns with the conventional practice in the XAI domain, commonly called sensitivity analysis.
@@ 4.7.2 Permutation Feature importance
The permutation feature importance (PFI) scores indicate the degree to which each feature influences the predictive performance of MCDFN. Features with higher importance scores contribute more significantly to the accuracy of the demand forecasts generated by the model. The PFI score for a feature $f$ is calculated as follows:
$\text{PFI}(f)=\frac{\text{Performance}_{\text{original}}-\text{Performance}_{\text{permuted}}}{\text{Performance}_{\text{original}}}$ (28)
where, $\text{Performance}_{\text{original}}$: Model performance metric (e.g., MSE, MAE) on the original dataset, $\text{Performance}_{\text{permuted}}$: Model performance metric after permuting the values of feature $f$.
Figure 13 presents the PFI analysis for the MCDFN model. The results reveal that cyclic temporal features such as the “Week_sin” feature emerge as the most influential predictor, followed by “Month_cos” and “Week_cos,” indicating the pronounced impact of weekly and monthly cyclical patterns on demand fluctuations. Moreover, the temporal features related to year and month exhibit substantial importance, underscoring the seasonal trends inherent in demand dynamics.
@@ 5 Discussion
The MCDFN can seamlessly integrate into modern SC management platforms to enhance demand forecasting capabilities. Many enterprises rely on Enterprise Resource Planning (ERP) systems, inventory management tools, and decision support systems to optimize SC operations. MCDFN can be deployed as a plug-and-play predictive module that integrates with widely used ERP solutions such as SAP, Oracle SCM, and Microsoft Dynamics, as well as inventory management tools like NetSuite and Fishbowl Inventory. This integration can be achieved through API-based deployment, allowing MCDFN to retrieve historical sales, real-time transactions, and inventory data to enhance forecasting. Alternatively, MCDFN can be incorporated into a data pipeline, enabling a seamless flow from data ingestion to forecasting to automated stock adjustments. Forecast outputs can be automatically fed into inventory management modules to adjust stock levels dynamically, reducing manual intervention in procurement decisions. MCDFN’s explainability features, such as ShapTime and Permutation Feature Importance, also allow SC managers to interpret key demand drivers, making inventory restocking decisions more transparent.
@@
Demand forecasting in this study was performed at an aggregate product family level, capturing broader demand patterns before refining insights for operational use. Our results confirm the importance of macro-economic factors in aggregate demand prediction, with the MCDFN model effectively capturing and incorporating complex, non-linear relationships among multiple predictive variables. Improved forecasting accuracy offers significant economic benefits, including reduced stockouts and overstocking, lower inventory holding costs, reduced obsolescence risk, and improved working capital efficiency. These benefits extend to better production scheduling, optimized logistics, and reduced emergency procurement costs, all of which contribute to lower total SC costs and improved service levels. Accurate aggregate forecasts can also be translated into more precise SKU-level predictions, potentially reducing inefficiencies and the bullwhip effect prevalent in non-seasonal demand industries (Cachon et al.,, 2007).
Effective demand forecasting significantly enhances SC efficiency, but fully eliminating the bullwhip effect requires a system-wide perspective and improved collaboration among SC participants (Gaur et al.,, 2005). While SC integration remains challenging, leveraging big data analytics (BDA) capabilities can yield favorable outcomes, particularly for upstream firms with limited visibility into end-consumer demand. Improved demand visibility also fosters collaboration, encouraging joint planning and information sharing, which is critical to long-term SC resilience. Despite adopting sophisticated analytical methods, many firms continue to rely on judgmental adjustments in demand forecasting (Boulaksil and Franses,, 2009; Moritz et al.,, 2014; Sanders and Manrodt,, 2003). Our findings confirm the ongoing importance of expert judgment in refining model outputs. As highlighted by Sanders and Manrodt, judgmental adjustments remain common even among practitioners using model-based forecasts, underscoring the enduring value of domain expertise alongside advanced analytics (Sanders and Manrodt,, 2003).
It is challenging to directly compare the performance of our proposed system with existing studies due to the absence of research incorporating a similar combination of deep learning techniques, decision integration strategies, and diverse learning algorithms in the demand forecasting process. Another limitation is the lack of real-world datasets like ours, making it difficult to benchmark against state-of-the-art studies. While this study presents the results of our proposed system, we also report findings from several related research efforts within the demand forecasting domain. A greedy aggregation-decomposition method was used to solve an intermittent demand forecasting problem for a fashion retailer in Singapore. The approach combined aggregated and detailed forecasts, improving accuracy compared to standard methods. The study achieved a 5.9% MAPE using a small dataset, showing that the method works well for handling irregular demand (Li and Lim,, 2018). A deep learning-based demand forecasting model was developed and tested using real-world retail data from the SOK Market in Turkey. The study combined deep learning, support vector regression, and multiple time-series models with a novel decision integration strategy. The approach improved forecast accuracy compared to single models. The proposed system achieved a MAPE of 24.7%, significantly reducing prediction errors. The research highlights the benefits of integrating different forecasting techniques for better accuracy(Kilimci et al.,, 2019). Both studies’ datasets are not publicly available due to confidentiality agreements, making it difficult to validate the MCDFN model on different industries. Without access to diverse real-world datasets, generalizing the model’s effectiveness beyond the original domain remains challenging.
@@ 6 Conclusions
Despite the promising results, our study has several limitations. First, the computational complexity of MCDFN is higher than traditional models, which may limit its applicability in real-time forecasting scenarios or environments with limited computational resources. Second, the performance of MCDFN heavily relies on the quality and quantity of the input data, making it sensitive to noisy or incomplete datasets. Third, while XAI techniques provide some level of interpretability, the model’s complexity can still challenge stakeholders to understand the underlying decision-making processes fully. Fourth, our study considered a single product family and a retail shop, potentially introducing selective biases due to geographical positions and other demographic variables.
Future research could address these limitations by exploring several avenues. Conducting a sensitivity analysis of window size to assess its impact on model performance could help fine-tune MCDFN for different forecasting scenarios and improve its adaptability. Implementing advanced data augmentation techniques and robust preprocessing pipelines would enhance the model’s ability to handle noisy or incomplete datasets, ensuring more reliable and stable predictions. Since most relevant studies’ datasets are confidential and not publicly available, making it difficult to validate MCDFN across different industries, future work should focus on obtaining diverse datasets or using synthetic data to evaluate generalizability. Combining MCDFN with other state-of-the-art techniques, such as attention mechanisms and transformers, could improve its predictive capabilities and interoperability. Exploring the integration of MCDFN with optimization techniques like Genetic Algorithms, Particle Swarm Optimization, or Mixed-Integer Programming can enhance SC decision-making.
Additionally, reinforcement learning could be applied to dynamically adjust demand forecasts in response to real-time changes, improving inventory control and order fulfillment. Comparing performance with generalized DL forecasting methods developed by industry experts, such as N-BEATS (ElementAI) and DeepAR (Amazon), can benchmark MCDFN’s effectiveness. Investigating the scalability of MCDFN in distributed computing environments and deploying it in real-time systems can demonstrate its practical applicability. Extending MCDFN to other domains, such as healthcare or finance, can validate its versatility and generalizability across different types of time series data. Identifying auxiliary variables that affect product demand and studying their influence on forecasting performance, along with integrating external factors like macroeconomic indicators, weather patterns, and social media trends, could further enhance demand forecasting accuracy.